### v1

# Controllable Learning in Information Retrieval: A Survey

Abstract
Controllable learning (CL) emerges as a critical component in trustworthy machine learning, ensuring that learners meet predefined targets and can adaptively adjust without retraining according to the changes in those targets. We provide a formal definition of CL, and discuss its applications in information retrieval (IR) where information needs are often complex and dynamic. The survey categorizes CL according to who controls (users or platforms), what is controllable (e.g., retrieval objectives, users’ historical behaviors, controllable environmental adaptation), how control is implemented (e.g., rule-based method, Pareto optimization, Hypernetwork), and where to implement control (e.g.,pre-processing, in-processing, post-processing methods). Then, we identify challenges faced by CL across training, evaluation, task setting, and deployment in online environments. Additionally, we outline promising directions for CL in theoretical analysis, efficient computation, empowering large language models, application scenarios and evaluation frameworks in IR.
Index Terms:
Controllable learning, trustworthy machine learning, information retrieval
1 Introduction
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=base east, child anchor=base west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=13em,fill=fill_0,draw=draw_0,font=,, where level=1text width=9.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=11.4em,fill=fill_2,draw=draw_2,font=,, where level=3text width=12.4em,fill=fill_3,draw=draw_3,font=,, where level=4font=,, [ Controllable Learning (CL)
, minimum height=3em [ Formulation of CL
(Sec. 2) [ Definition of CL
(Sec. 2) [ Define a task requirement triplet, $\mathcal{T}=\{\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}\}$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ represents the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ represents the context related to the task, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ represents the task target. ,leaf, text width=31.8em ] ] ] [ Taxonomy of CL
(Sec. 3) [ Who controls
(Sec. 3.1) [ User-centric control
(Sec. 3.1.1) [ Prior profile modification which empowers users to actively shape their recommendation experience: UCRS [1], IFRQE [2], LACE [3], and  [4]. , leaf, text width=31.8em ] ] [ Platform-mediated control
(Sec. 3.1.2) [ Impose algorithmic adjustments and policy-based constraints on the recommendation process: ComiRec [5], CMR [6], CCDF [7] and  [8]. , leaf, text width=31.8em ] ] ] [ What is controllable
(Sec. 3.2) [ Multi-Objective Control
(Sec. 3.2.1) [ Taking several objectives (metrics) as control targets, these work includes ComiRec [5], UCRS [1], CMR [6] , leaf, text width=31.8em ] ] [ Historical Behavior Control
(Sec. 3.2.2) [ Taking historical behavior as control targets , these work includes LACE [3], LFRQE [2], Grapevine [4] , leaf, text width=31.8em ] ] [ Controllable Environmental
Adaptation(Sec. 3.2.3) [ Taking environment adaptation as target, these work includes HyperBndit[9] and so on. , leaf, text width=31.8em ] ] ] ] [ How control is
implemented (Sec. 4) [ CL Techniques
(Sec. 4.1) [ Rule-Based Control
(Sec. 4.1.1) [ Rule-based post-processing techniques: [10], [11], [12], [13], [14] and [15]. , leaf, text width=31.8em ] ] [ Pareto optimization
(Sec. 4.1.2) [ Achieving controllability by using Pareto Optimization: PHN-HVI [16], PHNs [17],  [18],  [19],  [20],  [21],  [22] and  [23]. , leaf, text width=31.8em ] ] [ Hypernetwork
(Sec. 4.1.3) [ Hypernetworks based methods: Hamur[24], CMR[6], Hyperprompt[25], HyperBandit[9], [26], [27], [16], and [17] , leaf, text width=31.8em ] ] ] [ Where to control
(Sec. 4.2) [ Pre-processing methods
(Sec. 4.2.1) [ Including data manipulation methods and feature engineering methods: UCRS [1], LACE [3],  [2] and MetaLens [28]. , leaf, text width=31.8em ] ] [ In-processing methods
(Sec. 4.2.2) [ Including regularization and constrained optimization methods:CMR [6], CCDF [7]. , leaf, text width=31.8em ] ] [ Post-processing methods
(Sec. 4.2.3) [ Including reranking and result diversification methods: ComiRec [5] and UCRS [1]. , leaf, text width=31.8em ] ] ] ] [ Evaluations for CL
in IR (Sec. 5) [ Metrics (Sec. 5.1), minimum height=3em[ Single-objective metrics
(Sec. 5.1.1) [ NDCG [29], Recall,  Precision,  Hit Rate,  $\alpha$-NDCG [30], ERR-IA [31], Coverage, Iso-Index. , leaf, text width=31.8em ] ] [ Multi-objective metrics
(Sec. 5.1.2) [ Typical Multi-objective metrics include 4 main groups [32]. However, only the Hypervolume indicator [33] is most suitable in the current discussion context. , leaf, text width=31.8em ] ] ] [ Datasets (Sec. 5.2), minimum height=3em[ Amazon [34, 35], Ali_Display_Ad_Click [36], UserBehavior [5], MovieLens, MS MARCO [37] , leaf, text width=31.8em ] ] ] [ Challenges in CL
for IR (Sec. 6)[ Balancing difficulty
(Sec. 6.1)[ Pursuing controllability often leads to a trade-off, potentially compromising performance or other user-centric optimization metrics and adversely impacting accuracy or user experience. ,leaf, text width=31.8em ] ] [ Absence of evaluation
(Sec. 6.2)[ The assortment of perspectives on controllability and the consequent need for tailored evaluation metrics prevents direct methodological comparisons and hinder the progression of the field. , leaf, text width=31.8em ] ] [ Descriptions setting
challenges (Sec. 6.3)[ The task description serves as the instructions given by humans to the learner. A crucial issue is how to set the task target and transform it into a human understandable and precise description. ,leaf, text width=31.8em ] ] [ Online challenges
(Sec. 6.4)[ Scalability in realworld IR systems, particularly those dealing with streaming data and requiring continuous learning, is a formidable challenge. , leaf, text width=31.8em ] ] ] [ Future directions of
CL (Sec. 7)[ Several categories
(Sec. 7)
Figure 1: Overview of the survey of controllable learning (CL). This survey includes formulation of CL, taxonomy of CL, implementation methods of CL, how to evaluate the effectiveness of CL in information retrieval (IR) applications, and challenges of applying CL in IR.
Trustworthiness in machine learning is becoming increasingly important, with considerations such as controllability, fairness, privacy, and interpretability gaining prominence [38, 39, 40]. In this paper, we focus on the controllability of machine learning methods and their applications. Norbert Wiener, the originator of cybernetics, foresaw the ethical challenge of learning machine over sixty years ago when he stated [41]As machines learn they may develop unforeseen strategies at rates that baffle their programmers.” Additionally, Wiener explained the significance of controllability in learning machines [41]: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot efficiently interfere once we have started it, because the action is so fast and irrevocable that we have not the data to intervene before the actions complete, then we had better be quite sure that the purpose put into the machine is the purpose which we really desire and not merely a colorful imitation of it, etc.”, emphasizing two fundamental aspects of AI controllability: (1) A controllable learning machine needs to ensure that its predictions/decisions meet the target of the AI user; (2) Intervening in a learning machine after its deployment is a crucial but challenging task.
Controllability has been extensively discussed in the field of generative machine learning models, also known as controllable generation. For instance, in text generation [42, 43, 44] or visual generation [45, 46, 47], the AI models take a given task description (often referred to as a “prompt”) as input and generates content that aligns with that description. However, in the realm of discriminative machine learning models and applications, there lacks a unified definition and in-depth discussion of controllability. The current popular prompt-based controllable generation methods can be seen as a form of pre-processing, where, under fixed model parameters, changing input features allows the model to meet the target of the AI user. Existing theoretical results suggest that controlling model parameters may lead to better convergence than controlling model inputs when given task requirements [27], revealing that achieving controllability in machine learning involves various methods beyond prompt-based approaches. Therefore, controllable learning (CL) is a crucial category worthy of comprehensive review and exploration in trustworthy machine learning, ensuring that learning models meet user expectations and adhere to ethical standards. Meanwhile, in the context of Information Retrieval (IR) applications, CL enables models to adapt to diverse task description (e.g, users’ information needs) at test time without extensive retraining, delivering personalized and relevant results. Given the above landscape, this survey attempts to provide a formal definition of controllable learning, classify and discuss its application paradigms in information retrieval. In summary, this survey focuses on the topic of controllable learning, which aims to find a learner capable of adapting to different task requirements without the need for retraining, thereby meeting the desired task targets of the AI users.
From a commercial perspective, the ongoing evolution of large-scale machine learning models, including advanced large language models [48] like ChatGPT, has prompted a transition in the deployment of machine learning (ML) algorithms, moving away from the conventional Software-as-a-Service (SaaS) [49] towards Model-as-a-Service (MaaS) [50]. MaaS provides users with access to trained ML models via user-friendly interfaces, such as application programming interfaces (APIs). More specifically, MaaS allows users to access the trained models by calling APIs without the need to undertake the significant costs associated with model training and updating. MaaS empowers both companies and individuals to leverage the powerful capabilities of large models. However, it also poses new challenges for the controllability of ML models. Specifically, it raises the question of how to make ML models recognize the requirements of different downstream tasks and output personalized results that align with task targets without the necessity to retrain the model.
In the context of MaaS, the demand for controllability of ML models in Information Retrieval (IR) applications becomes more pronounced. This is attributed to the inherent complexity of information needs in IR tasks, where precise and tailored results are critical. In traditional settings, such as keyword-based search engines, the retrieval process is relatively straightforward, relying on predefined rules and algorithms. However, with the advent of MaaS and the utilization of large language models, the landscape has evolved. Now, users expect finer-grained and context-aware responses based on their information needs. This heightened expectation emphasizes the importance of controllability of ML models. Specifically, there is a need for mechanisms that enable individuals or platforms to easily express and convey their specific requirements for IR tasks. For example, in an e-commerce setting, a user searching for “black T-shirt” may expect different results based on varying weights of evaluation metrics (such as favoring diversity to simultaneously expose clothing that pairs well with black T-shirts), or setting specific filters to exclude irrelevant content (such as not exposing previously purchased T-shirts). Achieving this level of controllability necessitates sophisticated ML models capable of dynamically adjusting model parameters or outputs based on task requirements, without extensive retraining. Therefore, in the realm of IR, MaaS not only provides unprecedented access to powerful ML models but also underscores the significance of controllability and customization in addressing the diverse needs of users across different domains and applications.
The current landscape of research and application deployment in the aforementioned areas motivates us to write a comprehensive survey on controllable learning. As shown in Figure 1, this survey will encompass various aspects including the formal definition of controllable learning, methods and applications in information retrieval, evaluations, challenges, and future directions.
  • Formal definition of controllable learning. The survey will begin by providing a clear and concise formal definition of controllable learning.
  • Applications of controllable learning in IR. We will then delve into the applications of controllable learning in the field of IR, exploring who controls, what is controllable, and how control is implemented in IR systems.
  • Future directions and challenges. Finally, the survey will discuss the future directions and potential challenges in the field of controllable learning.
Overall, the survey on controllable learning will serve as a valuable resource for students, researchers, practitioners, and policymakers interested in understanding the principles, applications, and future prospects of this important area of trustworthy machine learning.
1.1 Comparison with Existing Surveys
The survey on trustworthy recommender systems [51] also talks about controllability, which divides it into two distinct categories: explicit controllability and implicit controllability. The former allows users explicitly edit or update the user preferences, indicating the users are exactly aware of their favors. The latter means that users could indirectly fine-tune their preferences when dynamically interacting with the recommender systems through re-ranking, modifying historical information, and so on. Similarly, a survey on user control in recommender systems [52] also split the controllability into two parts, where one is control during the preference elicitation phase and the other is control in the context of recommendation results. These two parts is corresponding to explicit controllability and implicit controllability above.
The aforementioned surveys primarily focus on recommender systems where users either explicitly adjust their preferences or implicitly allow the system to infer preferences from their behavior. However, they lack a summary of how controllability can be formally defined in machine learning. Moreover, there may be other possibilities for controllability in a broader sense, such as platform-side controllability or controllability with multiple objectives, which are not thoroughly defined and categorized. These surveys are also relatively early, and many new controllable techniques (such as hypernetworks [53], large language models [48], etc.) have emerged since then. Considering the limitations of existing surveys, in this survey, we provide a formal definition of controllable learning, and discuss controllable learning and its applications in information retrieval.
2 Foundations of Controllable Learning
In this section, we first give a formal definition of controllable learning in terms of task requirements, and describe the procedure of controllable learning.
The concept of controllable learning is manifested across various AI research directions, yet there is currently lack of a unified formal definition. To explicitly express the ingredients of controllable learning and unify existing controllable learning methods, we first define controllable learning in terms of task specifications. In brief, controllable learning is the ability to find a learner that can adapt to different task requirements without the need for retraining, thereby meeting the desired task targets, as detailed in Definition 1.
Definition 1 (Controllable Learning).
Define a task requirement triplet $\mathcal{T}=\{\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}\}\in\Gamma$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ represents the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ represents the context related to the task, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ represents the task target. Given an input space $\mathcal{X}$ and an output space $\mathcal{Y}$, for a learner $f:\mathcal{X}\rightarrow\mathcal{Y}$, controllable learning aims to find a control function $h$ that maps the learner $f$, the task description $\bm{s}_{\mathrm{desc}}\in\mathcal{T}$, and the context $\bm{s}_{\mathrm{ctx}}\in\mathcal{T}$ to a new learner $f_{\mathcal{T}}$ that fulfills the task target $\bm{s}_{\mathrm{tgt}}\in\mathcal{T}$, i.e,
$f_{\mathcal{T}}=h(f,\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}}).$
The integration of the learner $f$ and the control function $h$ is called a controllable learner. Moreover, upon receiving a new task requirement $\mathcal{T}^{\prime}\in\Gamma$, the control function $h$ should be capable of outputting a new learner $f_{\mathcal{T}^{\prime}}$ without the need for model training, ensuring that $f_{\mathcal{T}^{\prime}}$ satisfies the task target $\bm{s}^{\prime}_{\mathrm{tgt}}\in\mathcal{T}^{\prime}$.
Unlike domain adaptation or transfer learning[54, 55, 56], controllable learning adjusts the model adaptively in the face of new task requirements, eliminating the need for retraining for new tasks during the deployment phase. More specifically, in the context of a controllable learner during the testing phase, adjustments based on the task description can be categorized as follows: altering model inputs qualifies as a pre-processing method, modifying model parameters constitutes an in-processing method, and changing model outputs can be regarded as a post-processing method. However, it also presents challenges during the training phase in controllable learning, particularly in guiding the control function to identify and adapt to various task requirements effectively.
For the task requirement in Definition 1, the task target $\bm{s}_{\mathrm{tgt}}$ can be considered as the ideal quantitative metric that the controllable learner aims to achieve. The task description $\bm{s}_{\mathrm{desc}}$ is the specific representation of $\bm{s}_{\mathrm{tgt}}$ that can be inputted into an algorithm. $\bm{s}_{\mathrm{desc}}$ can be represented as a natural language prompt or as a vector containing multiple weights, among other forms. The context $\bm{s}_{\mathrm{ctx}}$ in task requirements can consist of features like historical data and the current user’s profile, offering controllable learners more background knowledge for prediction.
Under the definition, the procedure of controllable learning could be shown as Figure 2. When the output space $\mathcal{Y}$ in Definition 1 is not labels or lists but instead generated content, it is referred to as controllable generation. In this survey, our primary focus lies on controllable learning methods within information retrieval applications, which typically belong to controllable discrimination rather than controllable generation.
Figure 2: The procedure of controllable learning in Definition 1.
3 Controllable Learning in Information Retrieval: Taxonomy
Controllable learning plays a crucial role in modern information retrieval (IR), where learning models must dynamically adapt to various task descriptions, such as users’ specific information needs, without requiring extensive retraining. This capability ensures the delivery of personalized and relevant search results, thereby enhancing user satisfaction in IR systems. Recent advancements underscore the importance of controllability in contemporary IR platforms [8, 1, 3]. The following sections delve into the taxonomy of controllable learning in IR, exploring who controls and what is controllable in IR systems. Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “who” and “what” categories. (Table I).
3.1 Who Controls
“Who controls” refers to who proposes the task description $\bm{s}_{\mathrm{desc}}$, whether it is the user or the platform. When users have different needs, such as wanting to protect their privacy, or the platform wishes to increase result diversity to ensure different items get exposure, a corresponding task description arises. This section will introduce existing work based on who proposes the task description to control the IR systems.
3.1.1 User-Centric Control
Systems like UCRS [1], IFRQE [2], and LACE [3] empower users to actively shape their recommendation experiences. UCRS allows users to issue control commands to mitigate filter bubbles [1], IFRQE lets users decide which of their interactions should contribute to the training of the recommendation model [2], and LACE enables users to edit concept-based profiles for personalized text recommendations [3]. These systems exemplify user-centric control by providing users with the tools to adjust and refine the recommendation process according to their preferences and needs. Additionally, the user-controlled recommender system for connecting students with advisors further emphasizes user-centric control by leveraging user feedback and selections to guide the recommendation process, ensuring that the recommendations are aligned with the user’s academic or research goals [4].
3.1.2 Platform-Mediated Control
Platform-mediated control is characterized by algorithmic adjustments and policy-based constraints that a platform can impose on the recommendation process. For instance, recent work aims to consider a general platform-side framework to control the algorithms in recommender systems and to avoid the creation of isolated communities and the filter bubble effect [8]. ComiRec captures multiple interests from user behavior sequences and generates the top N items controllable by the platform in the aggregation module [5]. The CMR framework employs a hypernetwork to dynamically generate parameters for re-ranking models based on different preference weights, allowing the platform to adapt recommendations to various user groups or environmental changes without requiring user intervention [6]. CCDF platform controls the diversity of categories in recommendations by selecting top categories based on user-category matching and constrained item matching, aiming to alleviate echo chamber effects and broaden user interests [7]. Huan et al. [57] introduces a groundbreaking approach to addressing the complexities of heterogeneity in MSR. By deploying the Scenario Aware Model-Agnostic Meta Distillation (SAMD) framework, the research offers a method for scenario-aware and model-agnostic knowledge dissemination across diverse contexts, enhancing the system’s ability to tailor recommendations precisely. Some studies introduce specialized networks to meet the controllable demands. DTRN [58] obtains task-specific bottom representation explicitly by making each task have its own representation learning network in the bottom representation modeling stage. HyperBandit [9] maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. These systems demonstrate how platform-mediated control can optimize recommendation strategies to meet specific objectives, such as increasing diversity or improving conversion rates, while maintaining the quality of recommendations.
3.2 What is Controllable
This section aims to clarify “what is controllable” for a controllable learner in information retrieval applications. Specifically, we explore the question: “What is task target $\bm{s}_{\mathrm{tgt}}$ in Definition 1 of controllable learning?” A task target emerges when preferences shift across multiple objectives, such as accuracy and diversity. Another task target involves ensuring that learners can adapt to their environment, given descriptions of various environmental elements. Additionally, the context $\bm{s}_{\mathrm{ctx}}$ also needs to be controllable. For example, certain sensitive information in $\bm{s}_{\mathrm{ctx}}$ needs editing, while the target requires ensuring that the controllable learner’s outputs remain consistent with a model trained based on the edited contexts. Following this classification, this section reviews existing research to identify which target can be effectively achieved.
3.2.1 Multi-Objective Control
Objectives like precision, recall, and diversity, are at the core of recommender system performance metrics. ComiRec [5] strikes a balance between accuracy and diversity through hyperparameters in the aggregation module by capturing multiple user interests and integrating them into the recommendation process. UCRS [1] introduces a system where users can proactively reduce filter bubbles, controlling the recommendations they receive via hyperparameters that affect accuracy and coverage. The CMR framework [6] uses a hypernetwork to dynamically adjust preference weights, aiming to balance accuracy, diversity, etc. through platform-defined preference weights. These systems collectively focus on achieving a harmonious balance among the various objectives of recommendation quality, ensuring that users receive accurate, diverse, and relevant recommendations.
3.2.2 Historical Behavior Control
Some work in recommender systems involves allowing users to manage the context including their past interactions and profile information. LACE [3] allows users to edit their profiles by selecting or deselecting concepts, directly influencing the recommendations. LFRQE [2] lets users delete or modify past interactions to refine their recommendation experience. Grapevine [4] makes personalization more transparent and user-controllable using “open user models”, making normally hidden user models visible to end users and allowing them to change their values, thereby affecting the results of personalization. Together, these methods empower users to shape their digital footprint and ensure that recommendations better align with their evolving interests and preferences.
3.2.3 Controllable Environmental Adaptation
Compared to dynamic online learning [59, 60, 61], Controllable environmental adaptation is an explicit method that involves describing environmental elements and inputting them into the controllable learner, rather than solely relying on environmental feedback to update the model. HyperBandit [9] utilizes a hypernetwork to explicitly model the binary relationship between periodic external environments and user interests. In the testing stage, HyperBandit takes time features as input and dynamically adjusts the recommendation model parameters to account for time-varying user preferences.
3.2.4 Other Targets
In addition to the aforementioned control targets, there are many other forms of task targets, such as Multi-Scenario Control[58] and Multi-Domain Control[24], depending on the application context of the CL algorithms.
4 How Control is Implemented
In Section 3, the existing works are categorized based on controllers and controllable objectives. In this chapter, we summarize the common control methods and their application positions), corresponding to the CL techniques and application position (i.e., where to control). Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “CL tech.” and “where” categories (see  Table I).
4.1 Controllable Learning Techniques
We analyze and summarize the CL techniques for implementation of the control function $h$, where $h$ maps the learner $f$ to a new learner $f_{\mathcal{T}}$ based on the triplet $\mathcal{T}$.
4.1.1 Rule-Based Techniques
Rule-based controllable techniques have proven to be indispensable in many scenarios such as recommender systems, information retrieval, and text generation due to its direct series mode after the upstream models. These methods, acting as a patchwork solution, enhance system performance such as security, fairness, and interpretability from a product perspective. Following Definition 1, the $s_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents the system performance we expect to achieve, such as security, fairness etc. The $s_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ is always implicit, embedded in the design rationale of pre-processing methods like user profiles modification and post-processing methods, such as the removal of offensive language in text generation and enhancing the diversity in information retrieval. Conversely, $s_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ always represents the output from upstream models, including recommendation lists, retrieved documents, and raw generated texts.
Rule-based techniques in information retrieval systems involves applying predefined rules to the outputs of AI models to refine and enhance the results. Parikh [10] discusses the use of augmented AI and computer vision to extract information from unstructured data. By employing rule-based methods, the system can ensure that the extracted information is accurate and relevant, improving the overall quality and reliability of the information retrieval process. Khosrobeigi et al. [11] present a rule-based approach to improve the performance of Persian Optical Character Recognition (OCR). By incorporating rule-based corrections, the system can significantly reduce errors and enhance the readability and correctness of the output text. Kang et al. [12] demonstrate the application of rule-based natural language processing (NLP) techniques to improve disease normalization in biomedical texts. This approach ensures that the generated text not only meets the syntactic requirements but also maintains semantic consistency and domain accuracy. Recommender systems also benefit greatly from rule-based post-processing to ensure fairness, diversity, and explainability of recommendations. Nandy et al. [13] explore how fairness can be achieved through post-processing in web-scale recommender systems. By applying rule-based adjustments, the system can mitigate biases and ensure that recommendations are fair and equitable across different user groups. Le et al. [14] combine embedding-based and semantic-based models for post-hoc explanations in recommender systems. Rule-based post-processing plays a crucial role in providing clear and understandable explanations for the recommendations, enhancing user trust and system transparency. Antikacioglu and Ravi  [15] focus on improving diversity in recommender systems through post-processing techniques. By incorporating rules that promote diversity, the system can offer a broader range of recommendations, preventing the issue of over-specialization and improving user satisfaction.
Rule-based techniques is powerful in making AI systems more controllable across various applications. It allows for the fine-tuning of AI outputs to meet specific requirements, ensuring accuracy, fairness, and reliability. As AI continues to evolve, the integration of rule-based post-processing methods will remain crucial in enhancing the controllability and overall performance of AI systems.
4.1.2 Pareto Optimization
In the domain of machine learning, the concept of controllability has garnered significant attention due to the necessity to balance multiple, often conflicting objectives inherent in lots kinds of tasks. The integration of Pareto optimality principles from multi-objective optimization (MOO) has proven particularly instrumental in crafting algorithms that can navigate these trade-offs effectively. Following Definition 1, the $s_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents a single Pareto optimal solution that simultaneously satisfies multiple target objectives. The $s_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ generally consists of a set of multi-objective weights or constraints that enable the system to achieve the target. The $s_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context.
Lin et al. [62] pioneers this field by proposing an algorithm that decomposes the MTL problem into a series of subproblems, each representing a unique trade-off preference. This approach yields a set of Pareto optimal solutions, allowing for flexibility in selecting the most appropriate solution based on specific task requirements. Furthermore, Lin et al. [19] advances this concept by introducing a framework that enables real-time adjustments to these trade-offs, thereby providing a more dynamic and adaptable system capable of responding to changing user preferences and task demands. In parallel, the quest for scalable solutions has led to the exploration of methods that approximate the Pareto front, a representation of the optimal trade-offs among objectives. Ruchte et al. [18] propose a technique that conditions deep neural networks directly on preference vectors, thereby generating a well-distributed set of Pareto solutions. This work addresses the computational challenges associated with deep learning and sets the stage for more efficient and effective exploration of the Pareto front in complex MTL scenarios. A novel framework PHN-HVI [16] is proposed, which uses a hypernetwork to generate diverse solutions and improve the quality of the Pareto front in multi-objective optimization problems by optimizing the Hypervolume indicator, significantly outperforming existing methods. Pareto HyperNetworks (PHNs) [17] is introduced by Navon et al., where a single hypernetwork approach is applied to efficiently learn the entire Pareto front for multi-objective optimization problems, generalizes to unseen operating points, and enables post-training selection of optimal models based on runtime preferences. Complementing these efforts, research has also focused on the continuous exploration of Pareto sets, which provides a more nuanced understanding of the trade-offs in MTL. Ma et al. [20] introduce an innovative method that constructs continuous approximations of local Pareto sets, facilitating the applicability of MOO principles to the realm of deep MTL. Some researches have encompassed multi-objective optimization and Pareto efficiency in recommendation algorithms [21][22] [23]. Lin et al. [21] introduces a general framework for reconciling objectives such as Gross Merchandise Volume (GMV) and Click-Through Rate (CTR) in E-Commerce Recommendation. This work is complemented by Ribeiro et al. [22], which proposes novel methods for aggregating and hybridizing existing recommendation algorithms to achieve a Pareto-efficient ranking. Similarly, a hybridization method for multi-objective in recommender systems is proposed by Ribeir et al. [23], which contributes to this discourse by employing an evolutionary search to combine algorithms that excel in different objectives, leading to a Pareto frontier that can be adjusted according to user needs.
4.1.3 Hypernetwork
In addition to implicit control methods such as retraining and fine-tuning used to regulate models, hypernetworks have gradually emerged as a key technique for enhancing model controllability due to their explicit control capabilities. The advent of hypernetworks has revolutionized the landscape of artificial intelligence, particularly in the realm of parameter control for various tasks such as information retrieval and recommender systems. Hypernetworks, which are neural networks designed to generate the parameters for another network, offer a flexible and efficient way to manage and adapt model parameters dynamically. Following Definition 1, the $s_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents an expected result output by the controllable learner in the specific domain or task. The $s_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ typically represents the description of the task or domain, usually serving as the input for the hypernetwork. The hypernetwork outputs the parameters for the controllable learner, enabling it to achieve the corresponding $s_{\mathrm{tgt}}$. The $s_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context.
Galanti and Wolf [27] discuss the modularity of hypernetworks, emphasizing their ability to decompose a complex task into smaller, manageable sub-tasks. This modular approach allows for the generation of highly specialized parameters tailored to specific retrieval tasks, thereby improving accuracy and efficiency. He et al. [25] introduce Hyperprompt, a prompt-based task-conditioning method for transformers that leverages hypernetworks. This approach dynamically generates prompts that condition the transformer model to adapt to different retrieval tasks, enabling more precise and context-aware information retrieval. Li et al. [24] propose Hamur, a hyper adapter designed for multi-domain recommendation. Hamur utilizes hypernetworks to adaptively generate domain-specific parameters, thereby enhancing the model’s ability to recommend items across different domains with high relevance and accuracy. Furthermore, Chen et al. [6] explore the use of policy hypernetworks for controllable multi-objective re-ranking in recommender systems. By leveraging hypernetworks to generate parameters that balance multiple objectives, such as relevance and diversity, the system can dynamically adjust its recommendations based on user preferences and system goals. Similarly, Shen et al. [9] propose HyperBandit, whcih maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. Mahabadi et al. [26] discuss parameter-efficient multi-task fine-tuning for transformers via shared hypernetworks. This method allows for the efficient sharing of parameters across multiple tasks, reducing the overall computational load and improving the scalability of transformer models. The shared hypernetwork generates task-specific parameters that fine-tune the model for each particular task, enhancing its adaptability and performance. In conclusion, hypernetworks represent a powerful tool for parameter control in information retrieval and recommender systems. By enabling dynamic and task-specific parameter generation, hypernetworks can significantly improve the performance and adaptability of AI models, paving the way for more sophisticated and responsive systems.
4.1.4 Other Methods
There are also some control methods in machine learning, which is not strictly belong to the above three classifications. Multi-VAE [63] provide the ‘knobs’ in a Disentanglement (Disent.) way for users, with each knob corresponding to an item aspect. The latent space where generative factors (here, a preference towards an item category like genre) are captured independently in their respective dimensions, thereby enabling predictable manipulations. CGIR [64] propose a weakly-supervised method that can learn a disentangled item representation from user-item interaction data and ground the semantic meaning of attributes to dimensions of the item representation. During inference, CGIR start from the reference item and “walk” along the direction of the modification in the item representation space to retrieve a sequence of items in a gradient manner. Test-Time Adaption (TTA) means the algorithms do not change the model through retraining manner. especially in the scenario of online recommendation, where time consuming is quite strict, some works focus on how to change the model to adapt the dynamic environment directly in test stage. Reinforcement Learning (RL) is a kind of classic method to achieve controllability. Since these algorithms learn from the interaction with the environment (actually the reward function of the environment), it is possible for us to control the algorithms through dedicatedly designing the reward function. Inspired by the exceptional general intelligence of Large Language Models (LLMs) [48], researchers have begun to explore their application in recommender systems and other fields. [65] construct Supervised Fine-Tuning (SFT) tasks, augmented with labels derived from a conventional recommender model, explicitly improving LLMs’ proficiency in adhering to recommendation specific instruction which could be regarded as task description $s_{desc}$ as defined in Sec.2.
Table I: Summary of representative controllable learning (CL) methods for information retrieval. “Who” denotes the entity responsible for control (see Section 3.1), including User-Centric Control (User) and Platform-Mediated Control (Platform). “What” refers to what is task target (see Section 3.2) including, Multi-Objective Control (MOC), Historical Behavior Control (HBC), Controllable Environmental Adaptation (CEA), Multi-Scenario Control (MSC), and Multi-Domain Control (MDC). “CL Tech.” describes the techniques for implementing control (see Section 4.1) including Rule-Based Techniques (RBT), Pareto Optimization (Pareto), Hypernetwork (HN), Disentanglement Latent Space (Disent.), Test- Time Adaption (TTA), Reinforcement Learning (RL), Supervised Fine- Tuning (SFT). “Where” indicates the position where control is applied in the testing stage (see Section 4.2) including Pre-Processing Methods (Pre.), In-Processing Methods (In.), and Post-Processing (Post.).
Method Information Paradigm of Controllable Learning
Method Year Who What CL Tech. Where
ComiRec [5] 2020 Platform MOC RBT Post.
Mult-VAE [63] 2021 User MOC Disent. In.
CGIR [64] 2021 User HBC Disent. Pre.
UCRS [1] 2022 User MOC RBT Pre.&Post.
LACE [3] 2023 User HBC Disent. Pre.
CMR [6] 2023 Platform MOC HN In.
SAMD [57] 2023 Platform MSC HN In.
DTRN [58] 2023 Platform MSC HN In.
BFB [8] 2023 Platform MOC RL In.
HyperBandit [9] 2023 Platform CEA HN&TDA In.
HAMUR [24] 2023 Platform MDC HN In.
PHN-HVI [16] 2023 Platform MOC HN&Pareto In.
CCDF [7] 2024 Platform MOC HN In.
IFRQE [2] 2024 User HBC TDA Pre.
RecLM-gen [65] 2024 User HBC SFT&RL In.
4.2 Where to Control in Information Retrieval Models
In this section, we categorize control learning methods into three distinct categories based on the inference process during the testing phase. Specifically, we classify them according to the time when the control function defined in Definition 1 affects the process: methods that adjust model inputs before inference are termed pre-processing methods, those that operate on model parameters during inference are termed in-processing methods, and those that act on model outputs after inference are termed post-processing methods.
4.2.1 Pre-Processing Methods
Pre-processing methods, also referred to as embedding methods [27], achieve the task target solely by transforming the model inputs without adjusting the model itself, given a task description. Transformation of model inputs includes directly concatenating the task description onto the original feature vector. These methods are akin to in-context learning methods in large language models, where the task description serves as the prompt.
Mysore et al. [3] introduce the LACE model, which crafts user profiles using intelligible concepts extracted from user-document interactions. This innovation permits users to modify their profiles, thereby directly influencing the recommendation outcomes. Through comprehensive offline assessments and user studies, LACE’s capacity to refine recommendation quality via user interaction has been convincingly validated. Lastly, Wang et al. [2] tackle the pivotal concerns of user privacy and the degree of control over data utilization in model training. This framework grants users the discretion to decide which of their interactions may be harnessed for training purposes, delicately balancing between optimizing recommendation effectiveness and honoring user preferences. The implementation of an influence function-based model, alongside an augmented model featuring multiple anchor actions, evidences the feasibility of reconciling high-quality recommendations with user consent. UCRS [1] imagines a counterfactual world where out-of-date user representations are discarded, and estimates their effects as the difference between factual and counterfactual worlds. After deducting such effects, incorporates the control command into recommender inference. As to user-feature controls, it revises the user feature specified by the control command (e.g., changing age from middle age to teenager) to conduct the final inference at the two levels. As to item-feature controls, UCRS adopts a user-controllable ranking policy to control the recommendations w.r.t. item category.
4.2.2 In-Processing Methods
In-processing methods adaptively adjust the parameters or hyperparameters of IR models upon receiving the task description and context to achieve the task target. The CMR framework [6] harnesses the input preference vector to guide the hypernetwork, which in turn generates network parameters tailored to the desired balance of objectives, thus achieving the desired control during the testing stage. HyperBandit [9] utilizes the periodic time information to inject to a hypernetwork, modeling the relationship between user preference with corresponding time block, achieving efficient user preference adaptation during testing stage. Similarly, the CCDF [7] employs a hyperparameter, denoted as ‘$k$’, to directly manipulate the number of categories presented in the top-$k$ recommendations, allowing for precise control over the diversity of the recommended content. These methods demonstrate the utility of in-processing strategies in enhancing the adaptability and performance of recommender systems.
4.2.3 Post-Processing Methods
Post-processing methods, including reranking and result diversification, are crucial for refining the output of recommender systems. ComiRec [5] leverages a final aggregation module that balances dual objectives—accuracy and diversity—through a weighted summation approach to determine the top-k recommendations. This method ensures that the final selection of recommendations is not only precise but also diverse, catering to a broader range of user interests. Similarly, the UCRS [1] employs a weighted summation technique that balances accuracy and coverage to mitigate filter bubbles, ensuring that users are exposed to a wider array of content beyond their established preferences. These strategies highlight the importance of post-processing in delivering a nuanced and balanced set of recommendations that align with both the system’s performance goals and the user’s discovery of diverse items. Some works achieve controllability by re-ranking. MMR [66] promote diversity of the recommendation lists generated by some models via re-ranking.
5 Evaluation for Controllable Learning in Information Retrieval Applications
As far as we know, controllable learning was first proposed and defined by us, although it has been implicitly used in many works. However, existing works still lack specific evaluation criteria for controllable learning, and datasets to verify controllability have also not been explicitly proposed. In this section, we provide some common metrics in IR and clarify that appropriate use of these metrics can verify controllability. We also present common datasets in IR.
5.1 Metrics
In the definition of controllable learning that we discussed, we expect the control function $h$ to output a new learner $f_{\mathcal{T}}$ that meets the task requirements $\mathcal{T}$. Therefore, in the evaluation phase, we need to assess whether $h$ can effectively control the output $f_{\mathcal{T}}$, specifically whether the performance of $f$ meets the task requirements. For example, using a parameter $\alpha$ to represent the degree of control over the performance $s$ of the output $f$ (such as NDCG, diversity, MAP, etc.), a simple case would be an approximately linear relationship between $\alpha$ and $s$ [5][6]. A straightforward idea is to calculate the correlation coefficient between $\alpha$ and $s$, such as the Pearson Correlation Coefficient [67] and the Spearman Rank Correlation Coefficient [68], to measure the control effectiveness of the control function $h$.
In specific fields like information retrieval, we can utilize the combined variation of multiple single-objective metrics to assess whether the performance of $f$ meets the task requirements. Aiming that, multi-objective optimization metrics are also necessary to assess that variation. In this section, we therefore introduce some common single-objective metrics used in the field of information retrieval and some common metrics on multi-objective optimization.
5.1.1 Single-Objective Metrics
NDCG [29] evaluates the effectiveness of information retrieval systems by accounting for both the relevance and rank of retrieved documents. It applies a logarithmic discount to lower-ranked items and normalizes against an ideal ranking for consistency. This makes NDCG essential for comparing search engine and recommender system performance across different queries.
Precision is a metric used to evaluate the accuracy of information retrieval systems by measuring the proportion of retrieved documents that are relevant. It is calculated as the number of relevant documents retrieved divided by the total number of documents retrieved. High precision indicates that the system retrieves mostly relevant items, minimizing the presence of irrelevant information. This metric is particularly important in scenarios where presenting irrelevant items can negatively impact user experience, such as in search engines, recommender systems, and spam filtering.
Recall measures the effectiveness of information retrieval systems by determining the proportion of relevant documents that have been successfully retrieved from the total relevant documents available. It is calculated as the number of relevant documents retrieved divided by the total number of relevant documents. High recall indicates that the system retrieves most of the relevant items, which is essential in tasks where missing relevant information can have significant consequences, such as medical research or legal discovery.
Hit Rate evaluates the performance of recommender systems by assessing the presence of relevant items within the top-N recommendations provided to users. It is calculated as the number of users for whom at least one relevant item is included in the top-N recommendations, divided by the total number of users. Hit Rate is particularly useful for understanding the effectiveness of recommendation algorithms in scenarios where presenting at least one relevant option can significantly impact user satisfaction, such as in e-commerce or content streaming platforms.
$\alpha$-NDCG [30] extends the NDCG metric to evaluate the diversity in information retrieval systems. By incorporating a parameter $\alpha$, it penalizes redundancy and rewards the retrieval of diverse, relevant documents. This makes $\alpha$-NDCG particularly useful for tasks like web search and recommender systems, where presenting varied content is crucial.
ERR-IA [31] evaluates the effectiveness of information retrieval systems by considering user intent in the evaluation process. ERR-IA extends the Expected Reciprocal Rank (ERR) metric by incorporating a probabilistic approach to user intents, ensuring that the evaluation reflects the variety of user needs. It models user satisfaction as a function of the relevance of retrieved documents and their alignment with multiple user intents, making ERR-IA particularly useful for web search and other scenarios where understanding and catering to diverse user intents is crucial.
Coverage evaluates the comprehensiveness of information retrieval systems by determining the proportion of relevant items retrieved out of the total relevant items available across different queries or datasets. It is calculated as the number of unique relevant items retrieved divided by the total number of unique relevant items. High coverage indicates that the system retrieves a broad set of relevant items, making this metric essential for applications where capturing a wide range of relevant information is important, such as in comprehensive research databases, recommender systems, and digital libraries.
Iso-Index is a metric used to assess the fairness and equity of information retrieval systems. It measures the isolation of certain groups within the retrieved results, indicating the extent to which specific groups are underrepresented or segregated in the search results. A lower Iso-Index value suggests less isolation and, therefore, a more equitable distribution of information across different groups. This metric is particularly important in contexts where diversity and fairness are critical, such as social media, job recommendations, and news aggregation.
5.1.2 Multi-Objective Optimization Metrics
When the task requirements involve multi-objective or multi-task scenarios, the optimization goals may conflict and have constraints among themselves. That is, the optimal solution is not a single point but a surface or curve. To measure the quality of the front generated by the control function, new evaluation metrics are required. However, due to the scarcity of work on controllable multi-objective optimization in the IR field, traditional multi-objective optimization evaluation metrics are rarely used in the IR domain. Here, we briefly introduce some evaluation metrics related to controllable multi-objective optimization.
According to [32], performance indicators in multiobjective optimization can be mainly divided into four groups: Cardinality indicators, Convergence indicators, Distribution and spread indicators, and Convergence and distribution indicators. However, in practical implementations, the Pareto front approximation is not obtained through evolutionary algorithms that yield a finite (discrete) set of points. For example, [5] achieves diversity control by adjusting the continuous value $\lambda$ in the aggregation module, and similarly, [6] controls diversity and accuracy by adjusting the continuous value $\lambda$ in the hypernetwork’s input weights. This results in a complete, continuous Pareto front approximation, making performance indicators like Cardinality indicators, which calculate the number of points in the Pareto front approximation, less suitable. Likewise, distribution and spread indicators, which calculate the dispersion among discrete points, have little practical significance.
Moreover, when discussing multi-objective optimization problems in the IR context, there is often no ground truth Pareto front for the problem, making it impossible to measure the proximity between the ground truth and the approximation. Finally, among the convergence and distribution indicators, only the hypervolume indicator [33] is more appropriate for the current discussion context, and its calculation is feasible. Below, we will introduce the hypervolume indicator.
The hypervolume indicator [33] is a widely used performance metric in multiobjective optimization that measures the volume of the objective space dominated by a Pareto front approximation relative to a reference point. This indicator provides a comprehensive assessment of both convergence and diversity of the Pareto front.
Mathematically, the hypervolume $\mathrm{HV}$ for a given Pareto front approximation $P$ and a reference point $r$ is defined as:
$\mathrm{HV}(P,r)=\\ \text{vol}\left(\bigcup_{x\in P}\left[f_{1}(x),r_{1}\right]\times\left[f_{2}(x),r_{2}\right]\times\cdots\times\left[f_{m}(x),r_{m}\right]\right)$ (1)
where $f_{i}(x)$ is the value of the $i$-th objective function for solution $x$.
The hypervolume indicator calculates the volume of the $m$-dimensional space covered by the union of hyperrectangles defined by each solution in $P$ and the reference point $r$. The higher the hypervolume value, the better the approximation of the Pareto front, indicating better convergence and spread.
The choice of the reference point $r$ is critical as it influences the hypervolume value. It is usually set to a point that is dominated by all solutions. The hypervolume indicator is a powerful tool for evaluating the quality of Pareto front approximations in multiobjective optimization, balancing both convergence to the true Pareto front and diversity of the solutions.
5.2 Dataset
In this section, we summarize the commonly used datasets for controllable learning in information retrieval applications. To meet the various control requirements mentioned in Section 3.2, the datasets need to include the corresponding features. For example, to control result diversity, the data needs to include category information of items; to protect user privacy or control user history, the data needs to include user profile and interaction history. Here, we summarize the commonly used publicly available datasets that can be used for controllable learning research:
Amazon [34, 35]: This dataset comprises 142.8 million product reviews from various categories on Amazon [^1] https://www.amazon.com/, along with user and item profiles. And it includes category information of items, which can be used for multi-objective control such as diversity and fairness. It also contains time information, allowing for the extraction of users’ historical sequences and subsequently controlling these historical sequences.
Ali_Display_Ad_Click [36] [^2] https://tianchi.aliyun.com/dataset/56: The dataset includes records for 1 million users and 26 million ad display/click logs, featuring 8 user profile attributes (such as ID, age, and occupation) and 6 item features (such as ID, campaign, and brand).
UserBehavior [5] [^3] https://tianchi.aliyun.com/dataset/649: It collects user behaviors from Taobao’s recommender systems [69]. This dataset includes all behaviors (such as clicks, purchases, add-to-cart actions, and likes) of approximately one million randomly selected users with activity between November 25, 2017, and December 3, 2017.
MovieLens [^4] https://grouplens.org/datasets/movielens: This dataset is a classical movie recommendation dataset. It includes dataset versions of various sizes, such as 100k, 1M, 10M, and 20M. It includes information on users’ gender, age, and occupation, as well as item category information.
MS MARCO [37] (Microsoft Machine Reading Comprehension): This extensive dataset is designed for evaluating machine reading comprehension, retrieval, and question-answering capabilities in web search scenarios. It includes two benchmarks: document ranking and passage ranking, encompassing a total of 3.2 million documents and 8.8 million passages. Compiled from real user queries extracted from Microsoft Bing’s search logs, each query is paired with annotated relevant documents. This dataset spans a wide variety of question types and document genres, aiming to assess the performance of GR systems in complex web search scenarios.
6 Challenges in Controllable Learning
In this section, we discussed the challenges that may arise when applying controllable learning methods in the domain of information retrieval.
6.1 Balancing Difficulty in Training
The difficulty of balancing controllability with performance and efficiency is the pivotal challenge. Pursuing controllability often leads to a trade-off, potentially compromising performance or other user-centric optimization metrics and adversely impacting accuracy or user experience. For instance, controllability may be sought through the manipulation of hyperparameters—like a balancing factor in a loss function or a direct evaluative indicator. For instance, ComiRec [5] adjusts the balancing factor within its aggregation module to enhance diversity. Empirical evidence suggests that while diversity improves, accuracy can be compromised to some degree.
6.2 Absence of Evaluation
The absence of standardized benchmarks and evaluation metrics also hinders the development of Controllable Learning for IR, likely attributed to the very beginning stage of controllability learning and the lack of consensus on such metrics. Despite the shared objective of amplifying recommendation diversity, methods like CCDF [7] and ComiRec [5] adopt disparate evaluative approaches—the former implictly measuring it by designing a specific scenario and leveraging Hit Ratio (HR), while the latter assesses the diversity of top N recommended items by analyzing inter-category difference. The assortment of perspectives on controllability and the consequent need for tailored evaluation metrics can prevent direct methodological comparisons and hinder the progression of the field.
6.3 Setting Task Descriptions in Controllable Learning
In the context of a CL framework, the task target determines what is controllable, while the task description serves as the instructions given by humans to the learner. A crucial issue is how to set the task target and transform it into a human-understandable and precise description. Task descriptions are not limited to vectors or text; they can also take the form of images, graphs, rules, and other formats.
6.4 Challenges in Online Environments
Scalability in real-world applications (e.g., IR systems), particularly those dealing with streaming data and requiring online learning, is a formidable challenge. While research on controllability has been extensive in offline environments, integrating these principles into streaming IR applications, like online learning and reinforcement learning (RL), is yet to be fully realized. For instance, Wang et al. has designed influence-function-based models tailored to user preferences [2], but these models aren’t equipped to handle the swift changes in preferences without undergoing a retraining process—impractical for the real-time demands of streaming applications. Consequently, there’s an imperative need for research to pivot more resources into online settings. Future work could prioritize the development of models that can adjust on-the-fly to changing data and preferences. This involves creating systems capable of incremental learning and employing real-time feedback to refine their performance continuously. Progress in this direction will be a significant step toward the practical deployment of controllable learning models in dynamic streaming IR environments.
7 Future Directions
Controllable learning is gaining increasing attention, even though much of the current focus remains within implicit domains. This section highlights several future development directions, based on the challenging issues and emerging technologies (e.g., large language models). It is hoped that this chapter will provide valuable inspiration for future research. Why should we control? Theoretical analyses of controllable learning. Finding the optimal hypothesis in the hypothesis space for controllable learning is a more challenging task compared to traditional machine learning. Specifically, given the task requirements, the essence of in-processing controllable learning methods lies in establishing a mapping between the task target and the model parameters. Due to the vast parameter space of current deep learning models, uncovering structural information within this space and understanding causal associations with the target require rigorous theoretical analysis and effective training methods. This stands as a critical future research direction in controllable learning.
Controllable decision-making models. In numerous practical applications such as IR, feedback often takes the form of bandit feedback (i.e., feedback on decisions not executed is unobserved). This presents new challenges for sequential decision-making models such as reinforcement learning. Balancing exploration and exploitation while achieving adaptive control over task requirements is a crucial issue in both the theoretical analysis and practical applications of controllable learning.
Empowering LLM-based AIGC through controllable learning. Existing controllable generation methods typically rely on large language models (LLMs), using natural language prompts to control the input and obtain controllable AI Generated Content (AIGC). However, exploring controllable learning techniques to manipulate model parameters or outputs for achieving more specific task targets (e.g., preferences across multiple objectives) remains an area requiring further investigation.
Cost-effective control learning mechanisms. As stated in the definition of controllable learning (Definition 1), compared to the original learner, a controllable learner requires additional assistance from control functions, which inevitably introduces additional computational cost. Given the substantial computational costs linked with large-scale models, investigating efficient and cost-effective control mechanisms becomes imperative for future research.
Controllable learning for multi-task switching. Currently, in the field of information retrieval, there is limited research on controllable learning specifically tailored for search. Most existing work focuses on recommender systems. Therefore, adapting and extending controllable learning methods to the search represents a key direction for future research. One possible direction is how to utilize the same controllable matching model to adaptively switch between search and recommendation tasks. More broadly, leveraging a small set of controllable learning models to address multi-task, multi-objective, and multi-scenario switching challenges will not only enhance flexibility in addressing varied task requirements but also drive the development of novel methodologies capable of adapting to dynamic environments.
Demand for resource and metrics. Despite its significance, controllable learning lacks dedicated datasets and standardized evaluation metrics. For instance, the collection or construction of labels or user feedback across multiple objectives or diverse task requirements is crucial for the training and testing of controllable learners. Addressing these gaps represents a pivotal area for future research.
8 Conclusion
The landscape of controllable learning (CL) has been significantly enriched through the integration of diverse methodologies aimed at enhancing the trustworthiness of machine learning. CL’s ability to dynamically adjust learner according to predefined targets and adapt without retraining when those targets evolve makes it a critical component in fostering reliable and adaptive machine learning models. Particularly within the context of information retrieval (IR), CL offers a means to address the complexity and dynamism inherent in information needs. However, there remains much to explore in terms of theoretical guarantees, computational efficiency, and the integration of LLMs into controllable learning. Further advancements in these areas could lead to more sophisticated, user-centric, and adaptable AI models that meet the diverse and evolving needs of IR applications.